1) Before the offer for Gillette was announced (i.e., at the close of trade on 1/26/2005), P&G’s stock price was $55.44, and the firm had 2,522.583 million shares outstanding (fully diluted). If the proposed merger produced ZERO synergy, what would happen to P&G’s stock price after the merger is completed?

  • Given: Gillette’s stock price at the close of trade on 1/26/2005 was $45, with 1,068.379 million shares outstanding and had an implied offer for its stock price at $54.05
  • Given the price of the Gillette acquisition is $57.177 billion, its share price at the close of trade 1/26/2005 is $45 and the firm has 1,068.379 million shares outstanding the exchange ratio offer for Gillette is 0.975.
Answer #1

a). P&G stock price after merger (cash offer):

Value of combined firm after merger = value of P&G + value of Gilette (since synergy is zero)

value of P&G = price per share*number of fully diluted shares = 55.44*2,522.583 = 139,852 million

value of Gilette = 45*1,068.379 = 48,077.06 million

Value of combined firm = 139,852 + 48,077.06 = 187,929.06 million

Amount offered for Gilette = implied offer per share*number of shares = 54.05*1,068.379 = 57,745.88 million

Value of P&G after the merger = value of combined firm - amount offered for Gilette

= 187,929.06 - 57,745.88 = 130,183.17 million

P&G price per share after meger = 130,183.17/2,522.583 = $51.61 per share

b). P&G stock price after merger (stock offer):

Exchange ratio is 0.975 so 1 share of Gilette gets 0.975 shares of P&G

Therefore, total shares of P&G given for total shares of Gilette are 1,068.379*0.975 = 1,041.670 million

So, total number of shares for P&G after the merger = old number of shares + new shares issued = 2,522.583 + 1,041.670 = 3,564.253 million

%age of the new firm held by Gilette shareholders = shares issued/total new shares of P&G = 1,041.670/3,564.253 = 29.23%

So, the acquisition cost of 57.177 billion is for 29.23% ownership of new firm.

100% value of new firm = 100%*57,177/29.23% = 195,641 million

P&G price per share after merger = 195,641/3,564.253 = $54.89 per share